OpenStudy (sosa4954):

If y=7.5 when x= 0.5 find y whenx=-0.3

OpenStudy (sosa4954):

@bohotness @autumnm @Cookiemonster32 @DemiUnicorn @EuroSpeeder @Godlovesme

OpenStudy (gottennis121):

Is there an equation to figure out?

OpenStudy (sosa4954):

yeah could you help me and take me step through step its slope and direct variation

OpenStudy (gottennis121):

sure!

OpenStudy (sosa4954):

well that is the equation

OpenStudy (anonymous):

if y=7.5 when x=0.5, find y when x= -0.3 --- Use a proportion: y/x = y/x --------- y/-0.3 = 7.5/0.5 ----- y = -0.3*15 --- y = -0.45 Correct me if I am wrong If im wrong I will Fan and Medal you If I am right I will still fan and Medal you

OpenStudy (sosa4954):

Its worth taking the risk thank you

OpenStudy (anonymous):

What do you think gottennis

OpenStudy (gottennis121):

:D I'm just not in that grade yet that has learned that stuff. :D I thought this type of equation:\[x \times y \div 4 = ? \] Something like that

OpenStudy (gottennis121):

I wanted to help, but I guess I really can't. sorry